OpenStudy (anonymous):

find the number of terms in this series, the first term, and the last term. then evaluate the sum.

OpenStudy (anonymous):

the equ\[\sum_{n=2}^{10}(\frac{ 1 }{ 2 }n+3)\]ation is

OpenStudy (anonymous):

help please @whpalmer4

OpenStudy (whpalmer4):

do you understand the notation?

OpenStudy (anonymous):

no i have never really done one of these

OpenStudy (whpalmer4):

Okay, you could read that as "the sum of the values of \(\frac{1}{2}n+3\) as \(n\) goes from \(2\text{ to }10\)"

OpenStudy (whpalmer4):

I don't think I'm really giving anything away by telling you what those values are: \[4,\frac{9}{2},5,\frac{11}{2},6,\frac{13}{2},7,\frac{15}{2},8\]

OpenStudy (anonymous):

so there are 9 terms and the 9 terms can be found using the equation \[\frac{ 1 }{ 2 }n+3\] right

OpenStudy (whpalmer4):

But did you calculate the sum?

OpenStudy (anonymous):

yes i did thank you i have another question for you im about to post it will you help

OpenStudy (whpalmer4):

sure. what did you get for the sum?